OpenStudy (anonymous):

Can some one help me do this integral using partial fractions (I know its simple but I'm just learning partial fractions) Integral of (x+2)/(x^2+4x)

OpenStudy (anonymous):

substitute: u=x^2+4x then du=2x+4=2(x+2)

OpenStudy (anonymous):

I know its a simple u-sub problem. But i would like to see it done with partial fractions

OpenStudy (anonymous):

then you'll have integral 1/2 du/u right? can you finish it?

OpenStudy (anonymous):

inik... I know how to do the integral.. But i would like to see it done with partial fractions... And i even KNOW how to do partial fractions but for B I keep getting 3/2's and idk what to do with a 3/2's exactly

OpenStudy (anonymous):

\[=1/2 \ln|x ^{2}+4x| + const\]

OpenStudy (anonymous):

oups... let me try

OpenStudy (anonymous):

A/(x+4)+B/x=(x+2)/(x^2+4x) Ax+B(x+4)=x+2 so A and B are both 1/2

OpenStudy (anonymous):

hope from here it is clear

OpenStudy (anonymous):

I set it up with A / x + B / (x-4) then for my B value I got 3/2's ... I guess I just don't know what to do after I get a fraction

OpenStudy (anonymous):

A/x + (Bx+C)/(x^2+4)

OpenStudy (anonymous):

I got A=1/2, B=-1/2 and C=1 so, 1/(2x) +(-x/2 +1)/(x^2+1) integrate

OpenStudy (anonymous):

\[=\int\limits_{ }^{}dx/(2x) +\int\limits_{ }^{}(1-x/2)/(x ^{2}+4)=...\] =\[1/2\int\limits_{ }^{}dx/x +\int\limits_{ }^{}(2-x)dx/2(x+4) =...\] should work!

OpenStudy (anonymous):

Tfraiz, I got you lol.

OpenStudy (anonymous):

You have \[\int\limits \frac{x+2}{x^2+4x}dx\] Start by factoring the x out of the bottom. \[\int\limits \frac{x+2}{x(x+4)}dx=\int\limits(\frac{A}{x}+\frac{B}{x+4})dx\] Getting a common denominator your NUMERATOR would be: A(x+4)+Bx=x+2 Try plugging x=0 then: A(0+4)+B(0)=0+2 Or 4A=2 A=(1/2) Then plug in x=-4 A(-4+4)+B(-4)=-4+2 or -4B=-2 B=1/2 So you have: \[\frac{1}{2}\int\limits \frac{dx}{x}+\frac{1}{2}\int\limits \frac{dx}{x+4}=\frac{1}{2}\ln|x|+\frac{1}{2}\ln|x+4|+C\]